OpenStudy (anonymous):

((Cos x)/(1+sinx)) + ((1+sinx)/(cosx)) = 2 sec x. Verify the identity

OpenStudy (anonymous):

@mathstudent55

OpenStudy (dumbcow):

combine fractions what do you get?

OpenStudy (anonymous):

2 cosx + 4 sinx= 2 sec x ?

OpenStudy (dumbcow):

no what is common denominator

OpenStudy (anonymous):

Im not sure. X

OpenStudy (dumbcow):

ok \[\frac{1}{x} + \frac{2}{x-1}\] how would you add these fractions?

OpenStudy (anonymous):

X(x-1)

OpenStudy (dumbcow):

good, and what would the numerator look like

OpenStudy (anonymous):

Or 1x+(2x-2)

OpenStudy (dumbcow):

not quite, you multiply numerators by missing factor so \[\frac{1}{x}+\frac{2}{x-1} = \frac{1(x-1) + 2(x)}{x(x-1)}\]

OpenStudy (anonymous):

Oh okày

OpenStudy (dumbcow):

now apply to trig fractions ... what is common denominator

OpenStudy (anonymous):

Okay please give me a second

OpenStudy (anonymous):

I dont see how will you get it by adding the fractions @dumbcow

OpenStudy (anonymous):

i would rather use \((1-a^2)=(1-a)(1+a)\) :)

OpenStudy (dumbcow):

@myko , you do after simplifying and using (sin^2 +cos^2=1) also the "1+sin" cancels out

OpenStudy (anonymous):

Cosx(cosx)+((1+sinx)(1+sinx)) / (1+sinx)(cosx)

OpenStudy (dumbcow):

very good

OpenStudy (dumbcow):

now multiply out the numerator and we will try to combine some terms remember to FOIL the (1+sin)(1+sin)

OpenStudy (anonymous):

Cosx^2 +(1+sinx)^2 /

OpenStudy (anonymous):

Isn't the denominator (1+sinx)(cosx)

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

ok but what is (1+sin)^2 ... multiply it out

OpenStudy (anonymous):

(1+sin)(1+sin)

OpenStudy (dumbcow):

FOIL or distribute (x+1)(x+3) = x^2 +4x+3

OpenStudy (anonymous):

1+ 2sin+ sin^2

OpenStudy (dumbcow):

yes good now there is the pythagorean identity you should have learned that says: sin^2 + cos^2 = 1

OpenStudy (anonymous):

I'm getting confused with so many steps :/

OpenStudy (dumbcow):

hmm ok i will write it out \[\frac{\cos^2 x + (1+\sin x)(1+\sin x)}{\cos x (1+\sin x)}\] \[= \frac{\cos^2 x + \sin^2 x + 2 \sin x + 1}{\cos x (1+\sin x)}\] \[= \frac{1 +2 \sin x +1}{\cos x(1+\sin x)}\] \[= \frac{2 +2 \sin x}{\cos x (1+\sin x)}\] \[= \frac{2(1 + \sin x)}{\cos x (1+\sin x)}\] \[= \frac{2}{\cos x}\] \[= 2 \sec x\]

OpenStudy (anonymous):

In the third step how did you get rid of cos?

OpenStudy (dumbcow):

by using the identity i mentioned sin^2 + cos^2 = 1 whenever you see "sin^2 + cos^2" you can replace it with "1"

OpenStudy (anonymous):

Oh okay I see

OpenStudy (anonymous):

I understand now that I see it in steps. So in the end I'm left with 2secx=2secx

OpenStudy (dumbcow):

yep, when verifying an identity once you show that right side = left side. you are done

OpenStudy (anonymous):

Oh okay and what if the sides do not equal?

OpenStudy (dumbcow):

then you have more work to do :) depending on the problem, it may be easier to start on right side or even change both sides so that you can show right side = left side

OpenStudy (anonymous):

Oh okay thanks soooo much for explaining it all in detail!

OpenStudy (dumbcow):

yw:)